How many gallons of a 70% antifreeze solution must be mixed with 70 gallons of 20% antifreeze to get a mix that is 60% antifreeze? I tried doing .70(x)+.20(x+70)=.6 and came up with a negative answer of -14.88. Please help
Answer by mananth(16946)

-------------- percent ---------------- Amount
Amtifreeze I 20 ---------------- 70 gal
Antifreeze II 70 ---------------- x gal
Mixture 60 ---------------- 70 + x gal

20 * 70 + 70 x = 60 ( 70 + x )
1400 + 70 x = 4200 + 60 x
70 x -60 x = -1400 + 4200
10 x = 2800
/ 10
x = 280 gal Antifreeze II

m.ananth@hotmail.ca